Ask your own question, for FREE!
Mathematics 12 Online
OpenStudy (anonymous):

If you roll one dice and it lands on a 6, what is the percent chance of you rolling a 6 on another roll?

OpenStudy (psymon):

Well, what you roll the first time doesnt chance the odds of what you roll thesecond time. the way its worded it almost seems like a trick question. The % chance of you rolling a 6 on any roll is the same, doesnt matter if you already rolled a 6. So the % chance would be 1/6 = 16.67%. Now if the problem was asking the percent change that you roll a 6 twice in a row thatd be different. But based on how its phrased, Id say its just 16.67%

Can't find your answer? Make a FREE account and ask your own questions, OR help others and earn volunteer hours!

Join our real-time social learning platform and learn together with your friends!
Latest Questions
Arriyanalol: i need help
1 day ago 7 Replies 1 Medal
Arriyanalol: CONGRATS to me
2 days ago 8 Replies 2 Medals
Legomaster346: What colleges are good that are near Arab Alabama?
1 week ago 5 Replies 0 Medals
Arriyanalol: i need help
1 week ago 6 Replies 2 Medals
Can't find your answer? Make a FREE account and ask your own questions, OR help others and earn volunteer hours!

Join our real-time social learning platform and learn together with your friends!